### v2

# Uncertainty-Aware Knowledge Tracing Models

Abstract
The main focus of research on Knowledge Tracing (KT) models is on model developments with the aim of improving predictive accuracy. Most of these models make the most incorrect predictions when students choose a distractor, leading to student errors going undetected. We present an approach to add new capabilities to KT models by capturing predictive uncertainty and demonstrate that a larger predictive uncertainty aligns with model incorrect predictions. We show that uncertainty in KT models is informative and that this signal is pedagogically useful. This is important for an educational learning platform where understanding student ability is necessary.
1 Introduction
Knowledge Tracing (KT) models are widely used to predict a student’s response to a future question given their interaction history. KT models have developed over time from earlier work on Bayesian KT models [10, 16], to deep learning KT models based on LSTMs [11], self-attention [9, 5] and graphs [14], to more recent LLM based KT models [12]. Despite advances in the underlying KT models, the performance difference in terms of accuracy is often relatively small [8] and KT models have low specificity, which results in incorrect answers going undetected [13].
In this work we address the issue of incorrect student responses going undetected by considering predictive uncertainty in KT models. We utilise the Monte Carlo Dropout approach [4] to estimate predictive uncertainty in four different KT models to compare the impact of predictive uncertainty. We show that the predictive uncertainty of all four models provides a meaningful signal to detect misclassified incorrect student responses, which is pedagogically important for any online learning platform to ensure that student misconceptions do not go undetected.
2 Related Works
Knowledge Tracing (KT) models the temporal dynamics of student learning by predicting knowledge states based on sequences of interactions with educational content [2]. Various machine learning methods, including deep neural networks, have been proposed to address this sequential prediction task.
Among these, deep sequential models adopt auto-regressive architectures to capture the progression of student interactions. Deep Knowledge Tracing (DKT) [11] was one of the first to use an LSTM to estimate students’ knowledge states over time. Subsequent work, such as KQN [6], enhanced KT by encoding student learning activities into knowledge state and skill vectors and modeling their interactions with the dot product.
Memory-augmented models incorporate external memory mechanisms to better capture latent relationships between knowledge concepts (KC) and student knowledge states. DKVMN [17] introduced a static key memory for KC representations and a dynamic value memory to track students’ evolving mastery levels.
Graph-based models use graph neural networks to model intrinsic relations among questions, KCs and question-response interactions. [7] enhanced the KT model by introducing a question-KC bipartite graph to capture intrinsic relationships at both the question and knowledge component levels, thereby augmenting their respective representations. [15] employed graph convolutional networks to capture interdependencies among KCs and questions more explicitly.
Attention-based models leverage attention mechanisms to dynamically weigh past interactions. SAKT [9] introduced a self-attention architecture to model the relevance of historical responses. AKT [5] introduced monotonic attention modules to model forgetting behavior.
Despite these advances, uncertainty modeling in KT remains underexplored. Most existing models produce point predictions without quantifying the confidence in knowledge estimates. This is an important limitation for real-world educational applications. UKT [1] is the only other work to the best of our knowledge that introduces a KT model to capture uncertainty in student learning by representing knowledge states as stochastic distributions and modeling their transitions with a Wasserstein self-attention mechanism.
While UKT provides an approach to modeling uncertainty, it introduces an entirely new architecture, which limits its applicability in real-world settings where existing KT models are already deployed. Adding uncertainty quantification on top of such models remains non-trivial with this method.
3 Predictive Uncertainty for Knowledge Tracing Models
3.1 Knowledge Tracing Models
We adapt the standard formulation of knowledge tracing [11], which focuses on predicting a student’s performance on the next exercise based on their time-series interaction history. Formally, for each student, the sequence of exercises is modeled as $\{e_{i}\}_{i=1}^{t}$, where each exercise $e_{i}$ is represented as a three-tuple: $e_{i}=\{q_{i},c_{i},r_{i}\}$. Here, $q_{i}$ denotes the question ID, $c_{i}$ represents the features associated with $q_{i}$, and $r_{i}\in\{A,B,C,D\}$ is the student’s categorical response corresponding to one of the multi-choice answer options. The features linked to the question may include knowledge component identifiers, question text, explanation text for the answer options, and other pedagogical attributes. The KT model $F_{\theta}$ aims to predict the categorical outcome $\hat{r}_{t+1}$ for the next exercise, conditioned on the history of prior interactions and the information of the upcoming question. The prediction is given by:
$\hat{r}_{t+1}=F_{\theta}\left(q_{t+1},c_{t+1},\{e_{i}\}_{i=1}^{t}\right)$
In this work, we extend the usual binary KT formulation to a multi-class setting. We predict not only whether a student will answer correctly, but also which specific distractor option they are likely to select in the case of an incorrect response. This adaptation enables a more fine-grained understanding of student misconceptions and provides deeper insight into learning behaviors in mathematics education.
3.1.1 Baseline KT Models
To benchmark KT approaches for mathematics education, we consider four models as instantiations of $F_{\theta}$: Deep Knowledge Tracing (DKT) [11], Self-Attentive Knowledge Tracing (SAKT) [9], Attentive Knowledge Tracing (AKT) [5], and our proposed LLM Transformer KT model. DKT models student learning with recurrent neural networks (RNNs). It takes as input the sequence of question identifiers and responses $(q_{i},r_{i})$ and, conditioned on the next question $q_{t+1}$, predicts the probability of a correct response $\hat{r}_{t+1}$ by capturing temporal dependencies in the interaction history. SAKT replaces recurrence with a self-attention mechanism, which allows the model to selectively attend to relevant past interactions when predicting performance on the next question. AKT extends this approach by introducing contextual attention that considers both the similarity between questions and an exponential decay to model student forgetting. Finally, our LLM Transformer KT leverages self-attention over concatenated embeddings, which are generated using the Qwen3 0.6B embedding model [18], of the question text and associated text-based question features, and applies cross-attention with the response embeddings.
3.1.2 Limitations of KT Models
A limitation of our chosen baseline KT models is that they are typically evaluated using point-estimate metrics such as accuracy, F1 or AUC, which quantify predictive performance but provide no measure of model uncertainty. Modeling uncertainty is pedagogically valuable in an online learning platform. In intelligent tutoring systems, cases of high uncertainty could be handled differently. For example, by deferring the decision to a human expert or by providing the student with additional questions rather than acting on a potentially erroneous prediction. To support this kind of uncertainty-aware reasoning, we propose a Bayesian Neural Network (BNN) formulation of KT based on Monte Carlo Dropout (MC Dropout). This enables explicit quantification of uncertainty while preserving the flexibility of neural architectures.
3.2 MCDropout for KT Uncertainty
We estimate predictive uncertainty using Monte Carlo Dropout (MC Dropout) [4]. At inference, dropout remains active and predictions are averaged over $M$ stochastic forward passes. Following the information-theoretic approach based on Shannon entropy [3], we quantify total uncertainty by the entropy of the aggregated predictive distribution:
$H\big(p(r_{t+1}\mid q_{t+1},\mathcal{H}_{t})\big)=-\sum_{k}\bar{p}_{k}\log\bar{p}_{k},\quad\bar{p}_{k}\;=\;\tfrac{1}{M}\sum_{m=1}^{M}p_{k}^{(m)}.$
where $\bar{p}_{k}$ denotes the mean predictive probability of class $k$ estimated via MCDropout. Here, $\mathcal{H}_{t}=\{e_{i}\}_{i=1}^{t}$ represents the interaction history of a student up to time $t$, with each exercise $e_{i}=(q_{i},c_{i},r_{i})$ defined by the question identifier $q_{i}$, its associated features or knowledge components $c_{i}$, and the observed response $r_{i}$. Higher entropy corresponds to greater uncertainty in the model’s prediction, whereas lower entropy reflects more confident predictions. In addition to predictive entropy, we compute the standard deviation of the model’s predictions across Monte Carlo samples as a second measure of uncertainty, where further details are provided in Section 0.A.3.
4 Experiments and Results
We have implemented the proposed approach for four different models. Three of these, DKT [11], SAKT [9], and AKT [5], are common baselines often seen in the literature, and the fourth is our proposed self-attention variant that uses an LLM embedding model to process question text rather than a learned embedding of question ids that is more commonly used.
The dataset used to train each of the models contains student responses to questions. Here, we have limited the student response history to contain 100 questions-answer pairs per student. Students answer questions in quiz sessions, where a student will answer 5 questions in one session, which generally increase in difficulty. Therefore, the 100 question-answer pairs will typically be a result of a student working on 20 quizzes. We provide further details on the datasets in Section 0.A.1.
Firstly, we look at what the uncertainty in KT models reveals and why this is useful for a model designed to predict future responses of students. Secondly, we compare the different models to see how each captures uncertainty.
4.1 Uncertainty Prediction in Knowledge Tracing Models
The first result that we show, in Figure 1, is that the mean total entropy of the model predictions is larger when the model prediction is incorrect. This is very useful because the measure of uncertainty of the model predictions shows that the model is most often uncertain when the model is incorrect. Having a model say when it is uncertain could be used in a downstream application to reduce the amount of incorrect results being used; for example, in a learning platform instead of making an incorrect prediction about a student’s ability on a question, this task could be deferred to an expert.
Figure 1: Total entropy distribution of each model’s predictions as a function of their correctness. Misclassified data points clearly demonstrate a higher level of uncertainty.
Figure 2: Box plot of total entropy of each model’s predictions split into whether the student chose a correct response or an incorrect response.
Figure 3: Box plot of each model’s prediction standard deviation split into whether the student chose a correct response or an incorrect response.
We also show in Figure 3 that the model predictions have a larger mean entropy when the student chooses an incorrect response. This is largely a result of the models more frequently making incorrect predictions when students choose an incorrect response [13]. We speculate that the models likely have higher entropy for predictions when students choose an incorrect response due to incorrect responses being a minority class in the dataset and also identifying the specific distractor a student chose is more difficult than deducing if they choose a correct response. This result is further confirmed in Figure 3 where we compare the standard deviation of predictions from each model. In Figure 3 the mean standard deviation of predictions from each model is larger when the student has chosen an incorrect response.
By utilising the uncertainty from the model we can avoid learning gaps going undetected by only making use of the model predictions that have a higher certainty.
4.2 How Do Different Models Capture Uncertainty?
Next, we compare the different models to see what differences in uncertainty they capture. Figure 3 shows that the mean total entropy is quite similar across each of the models. This means that the level of uncertainty in the final aggregated prediction is similar across each model. On the other hand, Figure 3 shows that by looking at the standard deviation over the model predictions there is more variation between different models. By looking at the standard deviation over the model predictions we can see that some models have larger disagreement between individual predictions. This shows that DKT has the largest mean model uncertainty, whilst our model using LLM embeddings has the lowest mean model uncertainty.
Figure 4: Comparison of each models predictive total entropy against the question number currently being predicted.
Figure 5: Comparison of each models prediction standard deviation against the question number currently being predicted.
We also analyse how uncertainty varies as the student has answered more questions giving the model access to more previous responses. Figure 5 shows how the total entropy varies with question number. Here we can see that DKT has a much higher entropy for the first question than all the other methods. This implies that attention based models become more confident in their initial prediction, which may suggest they can more easily learn a question bias from the training data than the LSTM based models. We also observe a repeating spiking pattern in the entropy value over the question number. This is a result of students answering questions in chunks of five questions per quiz session, which generally increase in difficulty. As a result the more difficult question generates more incorrect responses from students, which we have shown to produce a larger entropy value from the model. We show that this spiking behaviour strongly correlates with the question difficulty in Section 0.C.
Figure 5 also shows a difference between attention based models and the LSTM based model. The standard deviation of predictions from the LSTM based model tends to increase as the number of previous questions seen increases up to approximately 20 questions at which point it remains constant. Conversely, for the attention based models the standard deviation of model predictions tends to decrease as the model has seen more answered questions, indicating that the model is becoming more confident in its predictions over time. We also see that our LLM embedding based model has lower uncertainty than other models, especially for lower numbers of previously seen questions.
5 Conclusion
Our results show that uncertainty in KT models is informative and pedagogically useful for mathematics education. Higher uncertainty reliably co-occurs with incorrect predictions, flagging when the model is least trustworthy. This can trigger concrete actions: a dialogue-based diagnosis or route the learner to targeted practice on prerequisite concepts. Uncertainty can also guide personalised learning pathways in maths: lower difficulty when uncertainty spikes, and raise difficulty as confidence grows. While total predictive entropy is similar across models, the standard deviation across Monte Carlo predictions separates them. DKT shows higher variance, whereas our LLM-based model shows consistently lower predictive uncertainty, suggesting greater confidence and stability. Together, these findings motivate integrating uncertainty into KT pipelines. They also indicate that attention-based models with rich text representations provide more calibrated and actionable signals for instructional decisions in mathematics.
{credits}
[…]
The authors have no competing interests to declare that are relevant to the content of this article.
Appendix
Appendix 0.A Experimental Setup
0.A.1 Dataset Details
We use a filtered subset of question-response logs from Eedi, an online mathematics learning platform. An interaction is a single question–response pair.
The corpus contains 4,2574{,}257 unique questions. Each question is multi-choice, and has question-ID and question-text associated with it. For each question, there are four options to choose from: {A, B, C, D}. We have the labels for which of those options is correct, and the other three options are carefully curated distractors tied to a mathematics misconception. The additional features a question contains includes construct-text, construct-ID, explanation-text and misconception-text. A construct-text is the most granular level of knowledge related to question, for example, ‘Construct 1: Find 10 less than a given number’. A misconception-text describes a cognitive error per distractor option which leads to mistakes. It is generic, unlike an explanation-text per option, which tends to be specific to the question.
We first filter students to those with at least 100 recorded responses and then retain the final 100 responses per student to standardize sequence length. The resulting training split contains  11,994\,11{,}994\, students (average $100$ responses; 1,199,2661{,}199{,}266 total interactions), and the validation split contains  1,493\,1{,}493\, students (average $100$ responses; 149,278149{,}278 total interactions) as shown in Table 1. All splits are student-disjoint: validation students do not appear in training. During evaluation, models predict each of the $100$ student responses in sequence with access only to prior interactions (causal/left-to-right), never the current or future response.
Table 1: Dataset statistics after filtering. Interactions are question–response pairs. Splits are disjoint by student.
Split # Students Avg. responses Total interactions
Train 11,994 100 1,199,266
Val 1,493 100 149,278
0.A.2 Model Training
We train four knowledge tracing models: DKT [11], SAKT [9], AKT [5], and our LLM-based Transformer, under a unified setup. Model sizes are summarized in Table˜2. AKT is the largest (3.3  M parameters), while our LLM Transformer remains compact (2.1  M). Unless otherwise noted, all models share the global training configuration in Table˜3: Adam optimizer, learning rate $3\times 10^{-4}$, cross-entropy loss, batch size 64, and 100 epochs, with a linear warmup followed by a cosine scheduler. Model-specific hyperparameters are listed in Table˜4. We fix embedding and hidden dimensions to 128 and use one layer across models, with dropout tuned per architecture (e.g., 0.5 for SAKT, 0.2 otherwise). The LLM-based model consumes text embeddings truncated to 1024 dimensions, while AKT uses the kq_same setting. All other details are held constant to enable a fair comparison across architectures.
Table 2: Parameter counts and estimated memory footprint for each model. We report total trainable parameters (in millions) and the estimated model size (in MB).
Model Total params (M) Est. size (MB)
LLM Transformer (ours) 2.1 8.436
DKT [11] 1.2 4.890
SAKT [9] 1.7 6.990
AKT [5] 3.3 13.020
Table 3: Global training hyper-parameters used across all models.
Setting Value
Learning rate (LR) 0.0003
Optimizer Adam
Loss Cross-Entropy
Batch size 64
Epochs 100
Scheduler Linear (warmup), cosine (training)
Table 4: Model-specific hyper-parameters.
LLM Transformer DKT SAKT AKT
Dropout 0.2 0.2 0.5 0.2
Embedding dim 128 128 128 128
Hidden dim 128 128 128 128
LLM truncation dim 1024 - - -
kq_same - - - True
Num layers 1 1 1 1
0.A.3 Model Evaluation
We use standard deviation of the model’s predictions over Monte-Carlo samples as a metric for predictive uncertainty. The standard deviation per question per student is calculated as:
$\mathrm{\sigma}\big(p\big)=\mathbb{E}_{k}\big[\sigma(p_{k})\big],\quad\mathrm{\sigma}\big(p_{k}\big)=\sqrt{\tfrac{1}{M}\sum_{m=1}^{M}\left[\big(p_{k}^{(m)}-\mu\big)^{2}\right]}$
where $k$ denotes the class, $p_{k}^{(m)}$ denotes the predicted probability of sample $m$, $\mu$ is the mean predicted probability, and $\mathrm{\sigma}\big(p_{k}\big)$ is the standard deviation of the model’s predictions over $M$ Monte-Carlo samples.
Appendix 0.B Model Quantitative Predictive Accuracy
On the validation split for the multi-class KT task, all models achieve similar accuracy Table˜5. Validation accuracy across training epochs is shown in Figure˜6.
Figure 6: Validation Accuracy of MCDropout Models: LLM Transformer, DKT, SAKT and AKT
Our LLM Transformer attains the highest accuracy ($67.17\%$), narrowly ahead of DKT ($67.07\%$), AKT ($67.13\%$), and SAKT ($67.05\%$). F1 varies more across models: AKT yields the strongest F1 ($35.97\%$), followed by SAKT ($34.15\%$), our model ($33.90\%$), and DKT ($32.57\%$). We also report AUC as a threshold-independent measure of model discrimination. The AUC scores are tightly clustered ranging from 75.35% (DKT) to 76.56% (LLM Transformer). These results suggest that while overall correctness is comparable across architectures, we need to also consider other metrics for AI mathematics tutoring applications.
Table 5: Quantitative results across models for validation split for multi-class KT formulation.
Model Accuracy F1 AUC
LLM Transformer (ours) 67.17 33.90 76.56
DKT [11] 67.07 32.57 75.35
SAKT [9] 67.05 34.15 76.22
AKT [5] 67.13 35.97 76.48
Appendix 0.C Uncertainty Analysis
Figure 8 shows the mean question difficulty against each question number the student is answering. The mean question difficulty at each question position follows a spiking pattern due to students answering questions in quizzes of sets of five questions, where the questions generally increase in difficulty during the quiz. We also show that the predictive entropy and the question difficulty are strongly correlated in Figure 8. This explains why there is a larger spike in predicted entropy at the final question number in Figure 5.
Figure 7: Comparison of question difficulty against the question number that would be predicted by a model.
Figure 8: Correlation between the model predictive entropy and the question difficulty.